one of the three angles of a triangle is twice the smallest and another is thrice times the smallest. find the angles
Answer by Engg.N Magampon(6)

x-smallest
2x-2nd angle
3x-3rd angle
x+2x+3x=180
6x=180
6x/6 = 180/6
x=30
2x=2(30)
=60
3x=3(30)
=90
smallest angle or the 1st angle=30degrees
2nd angle = 60degrees
3rd angle = 90degrees